What are the mean and standard deviation of the data -6,12,2,-4,1,6,0,3?
a. the mean is 2 1/3 and the standard deviation is approximately 5.6252
b. The mean is 2 1/3 and the standard deviation is 5.
c. The mean is 1 3/4 and the standard deviation is approsimately 5.6252
d. The mean is 4.86 and the standard deviation is 5.63.

Answer by Cromlix(4381)

x =-6+ 12 + 2 + (-4) + 1 + 6 + 0 + 3 = 14
14/8 mean = 1 3/4
(x - mean)
-7.75, 10.25, 0.25, -5.75, -0.75, 4.25
- 1.75, 1.25.
(x - mean)^2
-60.0625 + 105.0625+ 0.0625+ 33.0625+ 0.5625+
18.0625 + 3.0625 + 1.5625
Total = 221.5
Divide by 8 - 1 = 7
= 31.643
Square root answer.
= 5.6252 (Standard deviation)
Answer c is correct
Mean is 1 3/4 and SD is 5.6252
Hope this helps :-)